OpenStudy (anonymous):

14) Find the slope and y-intercept of the line 4x+2y=12

OpenStudy (anonymous):

@realmadrid7

OpenStudy (anonymous):

Hiya! Let's try an easier question first: if I gave you the line y = 2x + 3, would you be able to giv eme the slope and y-intercept?

OpenStudy (anonymous):

(If you're not sure, just say so!) :-)

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Okay, basically, any line with the form \[y = mx + c\] will have a slope "m" and a y-intercept "c". So in the example I gave you, the slope would be 2, and the y-intercept would be 3. Does that make sense, so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Great! In that case, try this one, before moving onto your original question. Can you find the slope and y-intercept of y=5x+7 ?

OpenStudy (anonymous):

ummmm

OpenStudy (anonymous):

Clue from before: If you have y=mx+c, then the slope is m and the y-intercept is c.

OpenStudy (anonymous):

can you please show me how to do it the teacher is timin us ... then will you teach me

OpenStudy (anonymous):

I'm trying to help you :-)

OpenStudy (anonymous):

is the slope 12

OpenStudy (anonymous):

Not quite! Let me give you the example from before: If you have y = 2x + 3 then the slope is 2 and the y-intercept is 3.

OpenStudy (anonymous):

So if you have y = 5x + 7, what are the slope and the y-intercept?

OpenStudy (anonymous):

5 is the slope and 7 is the why intercept

OpenStudy (anonymous):

Yep! Well done :-) How about y = 10x - 5 ?

OpenStudy (anonymous):

10 is the slope and 5 is the y intercept

OpenStudy (anonymous):

Almost! But there's a minus sign! What do you think the effect of that might be?

OpenStudy (anonymous):

-5

OpenStudy (anonymous):

That's it! :)

OpenStudy (anonymous):

Okay, now the important thing is to make sure your equation has the form y=mx+c.

OpenStudy (anonymous):

This means that:

OpenStudy (anonymous):

1) There's only one y, not two (e.g. 2y=mx+c would be bad!)

OpenStudy (anonymous):

2) The y is by itself (so x + y = 5 would be bad!)

OpenStudy (anonymous):

So if we take your original question (4x + 2y = 12), how do you think we might start to turn it into the form y=mx+c ?

OpenStudy (anonymous):

Are you still there?